Illustrate graphically the economic effects of an export subsidy of 15% if the world price is 90. Compute the producer surplus. Numbers 8-10 please.
Price of Jet (millions Quantity of jets demanded Quantity of jets supplied 140 120 110 100 90 80 70 60 50 40 1200 1000 900 800 700 600 500 400 300 200 100 150 200 250 300 350 400 450 500 600 20 Draw the market supply and demand curves. What are the equilibrium price and equilibrium quantity-AA oua 1. 40
B. Illustrate graphically the economic effects of an export subsidy of 15% if the world price is $90. Compute the producer surplus. The government decides to give a subsidy of $15 to the import competing industries. By how much producer surplus will increase. Illustrate your answer with a graph. 9. 10. List three arguments in favor of protectionist policies and discuss their shortcomings.
Answer #1

q1) In diagram the demand curve is downward sloping while the supply curve is upward sloping . The equilibrium occurs where demand = supply .This occurs when price = $60 and quantity demanded and supply = 400 units .

Paice Pt 130 (20 H0 loo bo 20 wai

q2)个 zup 130 (20 40 1 He Lemand 20- c. loo 200 300 4 0 Soo 600洳600 900 lo o no 1200 vd

In the diagram consumer surplus = area between the demand curve and equilibrium price = area ABE and

producer surplus = area BCE. = area between the supply curve and equilibrium price .

consumer surplus = area of triangle ABE= = 1/2 *base* height

= 1/2*400 *(140 - 60 )

= 1/2*400 *80

= 16000

producer surplus = area BCE

PRODUCER surplus = area of triangle BCE = 1/2 *base* height

= 1/2*400 *60

= 1/2*400 *60

= 12000

Q3) digram for ques 3 and 4 ce个 130 120 10 lo0 е»

With price ceiling of $40 , demand exceeds supply .At this price . demand = 500 units and supply = 200 units

deadweight loss = area BEFG

= 1/2* sum of base * height

= 1/2* (200 + 400 )*20

= 6000

q4 ) if world price now becomes $50 , then at this price , demand = 450 units and supply = 300 units .

consumer surplus = area ARU = 1/2* 450 * 90

= 20250

producer surplus = CRM= 1/2*300 * 50

= 7500

( answered first 4 parts as per policy )